OpenStudy (anonymous):

Please help!!! Find the value of k so that the line through the given points has the given slope. (2,-3) and (k,7); m=-2

OpenStudy (anonymous):

Thanks, but how did you go from -2=10/(k-2) to k-2=-5

OpenStudy (bradely):

-2=10/(k-2) Cross multiply -2(k-2)=10 k-2=10/-2 k-2=-5